OpenStudy (anonymous):

Use trig identities to transform the left side into the right: (tan(x) + cot(x)) / tan(x) = csc^2(x)

OpenStudy (raden):

first, (tan(x) + cot(x)) / tan(x) = tan(x)/tan(x) + cot(x)/tan(x) , agree ? @SeamN59

OpenStudy (anonymous):

Agreed

OpenStudy (raden):

oke, then tan(x)/tan(x) = 1 cot(x)/tan(x) = (1/tan(x))/tan(x) = 1/(tan^^2 (x)) so, we get (tan(x) + cot(x)) / tan(x) = tan(x)/tan(x) + cot(x)/tan(x) = 1 + 1/(tan^2 (x))

OpenStudy (raden):

opsss. that should there is once ^ not ^^, LOL

OpenStudy (raden):

did you get it , @SeamN59 ?